# Data Privacy Vocabulary (DPV) — Version 2.0

### v2

@@ 6.2 Ontology Engineering Process
The DPVCG consists of experts from multiple disciplines — computer science, law, sociology, authorities, and others. The primary role of its members is to discuss and reach consensus on the scope and information to be represented in DPV. Ontology engineers are then responsible for providing the appropriate modelling of concepts and organisation of DPV as a semantic web resource. While the DPVCG did not formally establish an ontology engineering methodology, practices that were adopted and evolved in the community reflect commonly utilised engineering methodologies of NeOn [suarez2011neon] and LOT [poveda2022lot].
The development process generally contained a member of the group proposing addition or modification of a concept, with information shared potentially via the mailing list and/or GitHub repo [^19] https://github.com/w3c/dpv/issues, and discussions and decisions within the meetings. Domain experts offered their advice on the information being modelled and what aspects of this should be considered, which were then formalised and shared as proposals. The group then discussed and voted to resolve the proposal, with minutes of the meeting reflecting the discussions and resolutions.
@@
The data in spreadsheets was structured to support ontology/taxonomy creation while still being comprehensible to the non-semantic web members. For generating the RDF serialisations, a custom documentation generator was developed which downloaded the spreadsheets, extracted the concepts and serialised them into multiple RDF formats, and generated the corresponding HTMLs. Existing tools such as WIDOCO [garijo2017widoco] were not used due to their limitation in control of outputs — for example DPV has taxonomies of concepts for purposes, personal data, etc. which would be all grouped together in a giant list in WIDOCO outputs, and which in DPV outputs are grouped separately to support adopters seeing related concepts grouped together. WIDOCO also did not support the dual RDFS+SKOS and OWL modelling outputs of DPV, or the ReSpec template [^20] https://respec.org/docs/ common in W3C outputs.
DPV follows best practices and guidelines established within the semantic web community. Namely, the W3C Best Practices for Publishing Linked Data [^21] Best Practices for Publishing Linked Data (2014) https://www.w3.org/TR/ld-bp/, WIDOCO best practices [garijo2017widoco] with OOPS! [^22] https://oops.linkeddata.es/ and FOOPS! [^23] https://foops.linkeddata.es/ — DPV’s score has false negatives regarding use of SKOS annotations, this has been reported to the maintainers for evaluation, W3ID [^24] https://w3id.org/ for permanent IRIs, GitHub [^25] https://github.com/w3c/dpv for version control and collaboration, and Zenodo [^26] https://doi.org/10.5281/zenodo.12505840 for archival.
@@ 6.3 Implications of Using SKOS and OWL
In OWL2, all DPV concepts are defined as instances of owl:Class and associated with each other using rdfs:subClassOf. The use-case concepts can now be declared as either subclasses of DPV concepts or as instances. In either case, DPV concepts cannot be directly used with the property dpv:hasPurpose as they are not instances of dpv:Purpose (shown in red). If the use-case creates instances — as shown where ex:CampaignA is created as an instance of dpv:DirectMarketing — then the use-case concept can be correctly used with dpv:hasPurpose. However, when ex:CampaignB is later introduced, it cannot be subclass of ex:CampaignA as ‘subclass of instance’ is undefined [^27] It might be tempting to use punning here — but it would not be correct. Punning enables use of the same concept as class and instance in separate contexts whereas we have a concept as class and instance in the same context — which is undefined. in OWL2. The relation between the two use-case concepts thus must be defined using either SKOS (thereby mixing SKOS and OWL2) or by using or creating another property, e.g., dct:hasPart. At this point, the use-case should reengineer its ontology by creating a class representing ex:CampaignA and creating the necessary subclass and instances to represent the relationship with ex:CampaignB. This however does not solve the issue with directly using DPV concepts as instances.
Thus, the RDFS+SKOS model is suitable for when DPV is to be used as a controlled taxonomy with a ‘lightweight ontology’ that supports extending the taxonomy in use-cases. Conversely, the OWL2 model is better suited for cases where formal reasoning is needed, and where sufficient ontology engineering capabilities exist to address changes in use-cases. By providing both serialisations, DPV enables the adopters to choose the most suitable serialisation that supports their use-case and/or existing implementations, and retains semantic interoperability based on converting between SKOS and OWL2282828Using OWL and SKOS (2008) https://www.w3.org/2006/07/SWD/SKOS/skos-and-owl/master.html.
@@ 7 Conclusion & Future Work
The DPVCG plans to refine its TECH and AI extensions based on existing works [golpayeganiAIROOntologyRepresenting2022, golpayeganiAICardsApplied2024] providing taxonomies for AI techniques, capabilities, lifecycle stages, risks and risk sources, and to enable stakeholders to express specific use-cases (e.g., involving generative AI) in a manner that supports requirements for EU AI Act and ISO standards [golpayeganiAICardsApplied2024]. The DPVCG is also continuing its efforts to develop vocabularies to represent key ‘data and AI regulations’ notably in EU the Digital Services Act (DSA), Data Markets Act (DMA), Data Act, and Data Spaces, and modelling laws in other jurisdictions, e.g., Ireland, USA, and UK.
To support the application of DPV in regulatory environments, the DPVCG is developing guides based on existing work that utilises DPV to support GDPR implementations. These include Records of Processing Activities (ROPA) [ryan_dpcat_2022], Data Protection Impact Assessment (DPIA) [pandit_semantic_2022], Data Breaches [pandit_towards_2023], and Rights management. To support the implementation of decentralised and user-centric applications of DPV, such as those envisioned in the IDSA, IEEE P7012 and Solid, the DPVCG is developing vocabularies and guides — for example implementation of ISO Standards for (semantic) privacy notices. In addition, the DPV is also looking to incorporate the Standard Data Protection Model [^29] https://www.bfdi.bund.de/EN/Fachthemen/Inhalte/Technik/SDM.html provided by the German data protection authority regarding implementation of concrete technical and organisational measures.
@@
Acknowledgements of Funding
The DPVCG was established as part of the SPECIAL H2020 Project, which received funding from the European Union’s Horizon 2020 research and innovation programme under Grant#731601 from 2017 to 2019. Harshvardhan J. Pandit was funded to work on DPV from 2020 to 2022 by the Irish Research Council’s Government of Ireland Postdoctoral Fellowship Grant#GOIPD/2020/790. The ADAPT SFI Centre for Digital Media Technology is funded by Science Foundation Ireland through the SFI Research Centres Programme and is co-funded under the European Regional Development Fund (ERDF) through Grant#13/RC/2106 (2018 to 2020) and Grant#13/RC/2106_P2 (2021 onwards). The contributions of Piero Bonatti and Luigi Sauro to the DPVCG have been funded by the European Union’s Horizon 2020 research and innovation programme under grant agreement N. 731601 (SPECIAL) until 2019, and under Grant#883464 (TRAPEZE) from 2020 until 2023. The contributions of Beatriz Esteves and Delaram Golpayegani have received funding through the PROTECT ITN Project from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant#813497. Beatriz is also funded by SolidLab Vlaanderen (Flemish Government, EWI and RRF project VV023/10). The contributions of Julian Flake have received funding from the German Federal Ministry of Education and Research (BMBF) grant#16KIS1298 (AI-NET PROTECT), from the European Union’s Horizon Europe Framework Programme grant#101129822 (TITAN) and from the European Union’s Digital Europe Programme grant#101123471 (EDGE-Skills).
For the purpose of Open Access the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
[…]

### v3

@@ 6.2 Ontology Engineering Process
The DPVCG consists of experts from multiple disciplines — computer science, law, sociology, authorities, and others. The primary role of its members is to discuss and reach consensus on the scope and information to be represented in DPV. Ontology engineers are then responsible for providing the appropriate modelling of concepts and organisation of DPV as a semantic web resource. While the DPVCG did not formally establish an ontology engineering methodology, practices that were adopted and evolved in the community reflect commonly utilised engineering methodologies of NeOn [81] and LOT [69].
The development process generally contained a member of the group proposing addition or modification of a concept, with information shared potentially via the mailing list and/or GitHub repo, and discussions and decisions within the meetings. Domain experts offered their advice on the information being modelled and what aspects of this should be considered, which were then formalised and shared as proposals. The group then discussed and voted to resolve the proposal, with minutes of the meeting reflecting the discussions and resolutions.
@@
The data in spreadsheets was structured to support ontology/taxonomy creation while still being comprehensible to the non-semantic web members. For generating the RDF serialisations, a custom documentation generator was developed which downloaded the spreadsheets and serialised them into multiple RDF formats, and generated the corresponding HTMLs. Existing tools such as WIDOCO [35] were not used due to their limitation in control of outputs — for example DPV has multiple taxonomies which would be all grouped together in a giant list in WIDOCO outputs, and which in DPV outputs are grouped separately to support adopters seeing related concepts grouped together. WIDOCO also did not support the dual RDFS+SKOS and OWL modelling outputs of DPV, or the ReSpec template [^16] https://respec.org/docs/ common in W3C outputs.
DPV follows best practices and guidelines established within the semantic web community. Namely, the W3C Best Practices for Publishing Linked Data (2014) WIDOCO best practices [35] with OOPS! [^17] https://oops.linkeddata.es/ and FOOPS! [^18] https://foops.linkeddata.es/ for evaluation, W3ID [^19] https://w3id.org/ for permanent IRIs, GitHub for version control and collaboration, and Zenodo for archival.
@@ 6.3 Implications of Using SKOS and OWL
In OWL2, all DPV concepts are defined as instances of owl:Class and associated with each other using rdfs:subClassOf. The use-case concepts can now be declared as either subclasses of DPV concepts or as instances. In either case, DPV concepts cannot be directly used with the property dpv:hasPurpose as they are not instances of dpv:Purpose (shown in red). If the use-case creates instances — as shown where ex:CampaignA is created as an instance of dpv:DirectMarketing — then the use-case concept can be correctly used with dpv:hasPurpose. However, when ex:CampaignB is later introduced, it cannot be subclass of ex:CampaignA as ‘subclass of instance’ is undefined [^20] It might be tempting to use punning here — but it would not be correct as we have a concept as class and instance in the same context — which is undefined. in OWL2. The relation between the two use-case concepts thus must be defined using either SKOS (thereby mixing SKOS and OWL2) or using another property, e.g., dct:hasPart. At this point, the use-case should reengineer its ontology by creating a class representing ex:CampaignA and creating the necessary subclass and instances to represent the relationship with ex:CampaignB. This however does not solve the issue with directly using DPV concepts as instances.
Thus, the RDFS+SKOS model is suitable for when DPV is to be used as a controlled taxonomy with a ‘lightweight ontology’ that supports extending the taxonomy in use-cases. Conversely, the OWL2 model is better suited for cases where formal reasoning is needed, and where sufficient ontology engineering capabilities exist to address changes in use-cases. By providing both serialisations, DPV enables the adopters to choose the most suitable serialisation that supports their use-case and/or existing implementations, and retains semantic interoperability based on converting between SKOS and OWL2212121Using OWL and SKOS - Sean Bechhofer and Alistair Miles, W3C (2008)
@@ 7 Conclusion & Future Work
The DPVCG plans to refine its TECH and AI extensions based on existing works [37, 36] providing taxonomies for AI techniques, capabilities, lifecycle stages, risks and risk sources, and to enable stakeholders to express specific use-cases (e.g., involving generative AI) in a manner that supports requirements for EU AI Act and ISO standards [36]. The DPVCG is also continuing its efforts to develop vocabularies to represent key ‘data and AI regulations’ notably in EU the Digital Services Act (DSA), Data Markets Act (DMA), Data Act, and Data Spaces, and modelling laws in other jurisdictions, e.g., Ireland, USA, and UK.
To support the application of DPV in regulatory environments, the DPVCG is developing guides based on existing work that utilises DPV to support GDPR implementations. These include Records of Processing Activities (ROPA) [74], Data Protection Impact Assessment (DPIA) [60], Data Breaches [68], and Rights management. To support the implementation of decentralised and user-centric applications of DPV, such as those envisioned in the IDSA, IEEE P7012 and Solid, the DPVCG is developing vocabularies and guides — for example implementation of ISO Standards for (semantic) privacy notices. In addition, the DPV is also looking to incorporate the Standard Data Protection Model [^22] https://www.bfdi.bund.de/EN/Fachthemen/Inhalte/Technik/SDM.html provided by the German data protection authority regarding implementation of concrete technical and organisational measures.
@@
Acknowledgements of Funding The DPVCG was established as part of the SPECIAL H2020 Project funded by the European Union’s Horizon 2020 programme under Grant#731601 (2017-2019). Harshvardhan J. Pandit was funded (2020-2022) by the Irish Research Council’s Government of Ireland Postdoctoral Fellowship Grant#GOIPD/2020/790. The ADAPT SFI Centre for Digital Media Technology is funded by Science Foundation Ireland through the SFI Research Centres Programme and is co-funded under the European Regional Development Fund (ERDF) through Grant#13/RC/2106 (2018-2020) and Grant#13/RC/2106_P2 (2021 onwards). Piero Bonatti and Luigi Sauro were funded by the European Union’s Horizon 2020 programme under Grant#731601 (2017-2019), and under TRAPEZE Grant#883464 (2020-2023). Beatriz Esteves and Delaram Golpayegani were funded by European Union’s Horizon 2020 programme’s Marie Skłodowska-Curie Grant#813497 PROTECT ITN Project. Beatriz is also funded by SolidLab Vlaanderen (Flemish Government, EWI and RRF project VV023/10). Julian Flake received funding from the German Federal Ministry of Education and Research (BMBF) grant#16KIS1298 (AI-NET PROTECT), from the European Union’s Horizon Europe Framework Programme grant#101129822 (TITAN) and from the European Union’s Digital Europe Programme grant#101123471 (EDGE-Skills). For the purpose of Open Access the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript arising from this submission.